sarah bought 4 pounds of bok choy to make soup. if the recipe that calls for 7 ounces of bok choy for the following recipe, how many batches of the recipe can she make ?
Answer by Fombitz(32388)

4 lbs*(16 oz/1 lb)=64 oz of bok choy
N=64/7=9+1/7
She can make 9 batches and have a little left over.